OpenStudy (anonymous):

during a bicycle race, the racers must complete 210 miles. One cyclist, traveling 10 mph faster than the second cyclist, covers this distance in 2.4 hours less time than the second cyclist. Find the rate of the first cyclist. (Applications of Quadratic Equations)

OpenStudy (anonymous):

PLEASE HELP!

OpenStudy (bahrom7893):

uhmm, V1 = V2 + 10

OpenStudy (bahrom7893):

t1 = t2-2.4

OpenStudy (anonymous):

so how do i do that?

OpenStudy (bahrom7893):

D = V*t = (V2+10) * (t2-2.4) 210 = (V2+10) * (t2 - 2.4)

OpenStudy (bahrom7893):

was there anything else given?

OpenStudy (anonymous):

no, thats the whole problem

OpenStudy (bahrom7893):

no it wasn't... lol

OpenStudy (bahrom7893):

They have to cover the distance in ONE DAY

OpenStudy (bahrom7893):

nevermind, i still dont see how that helps...

OpenStudy (bahrom7893):

workin on it..

OpenStudy (anonymous):

okay, thanks

OpenStudy (bahrom7893):

what is rate, i think theyr not asking u to solve for velocity, they just want u to show how u can find it..

OpenStudy (bahrom7893):

wait do u have the answer?

OpenStudy (bahrom7893):

ughhh is this college algebra?

OpenStudy (bahrom7893):

Sorry i can't do anything else.. =( i did my best.. hate word problems!!!

OpenStudy (bahrom7893):

try yahoo answers maybe?